OpenStudy (anonymous):

On a coordinate plane, find the midpoint (M) of a segment with the following endpoints: A(-4, 13) and B(9, 0) 1) M(13, 5) 2) M(6.5, 2.5) 3) M(2.5, 6.5) 4) M(5, 13)

OpenStudy (anonymous):

x = (-4 + 9) / 2 = 5 / 2 = 2.5 y = (13 + 0) / 2 = 13 / 2 = 6.5 3 is your answer.